high fountain of water is located at the center of a circular pool not wishing to get his wet a student walk around the pool and measure its circumference to be 15.0 m. next the student at the edge of the pool and uses a protector to gauge the angle of alivation of the top of the mountain be 55.0 degree.How high is the fountain?
1
Expert's answer
2019-09-23T09:22:35-0400

First, lets find the radius of the circular pool. As the circumference is 15m:


"L=2\pi R \rightarrow R=\frac{L}{2\pi}=\frac{15}{2*3.14}=2.39m"

Then, the height of the fountain:


"H=R*\tan\theta=2.39*\tan55^{\circ}=3.4 m"

Answer: The height of the fountain is 3.4 meters